# Inga Strümke5, Vince I. Madai6,7, Tobias Budig8, Ali Sunyaev9, and Adam Hilbert10 On behalf of the VALIDATE consortium

### v1

@@ VI Discussion
A limitation of this study and suggestion for future work is that the used FL algorithm (e.g. FedAvg) was not extracted, which could be a relevant factor in analyzing the impact on explanations. As stated above, adherence to information reporting is important, and future reviews on the field should look closer at this aspect. Further, we conducted our search in 2023. In the ever-changing domain of ML model development, further developments could have been achieved in the meantime.
VII Conclusion
@@
This review highlights the interplay between FL and XAI, both methodologically and experimentally. We have identified a research gap in which few studies quantify the impact of FL on explanations. Moreover, the impact of FL on model interpretability and explanations remains unclear, revealing the need for studies that quantify such impact. There is a need for rigorous experimental and analytical research to assess how FL training influences the structure of the model and its implications for explainability and interpretability. Additional research should also provide guidance for practitioners on the responsible implementation of FL with XAI, particularly in critical fields such as healthcare, finance, and engineering. This guidance will help ensure responsible deployment of AI systems that balance model performance with due transparency.
@@
Many papers do not use consistent terminology for explainability, which complicated the analysis. It is important for future research to clearly define terms and use standardized nomenclature. Another important finding is that a minority of the studies specify which FL libraries were used. Future publications should include details on the libraries employed or provide source code for custom implementations. Furthermore, the lack of adequate reporting of data characteristics calls for strict adherence to reporting guidelines. These are important improvements towards ensuring reproducibility, transparency, and auditability, which are crucial ethical aspects in sectors such as medicine.
In conclusion, this review underscores the need for more structured and transparent research practices in the intersection of FL and XAI. Establishing clear definitions and consistent methodologies will be key in advancing the field and addressing the identified gaps. As demand for FL and XAI continues to grow, particularly in high-stakes environments, the importance of rigorous, transparent, and reproducible research cannot be overstated.
@@ A-A Combining FL and explanation methods: methodological contributions
[87] proposes the first counterfactual explanation method for VFL. They show the validity by retraining VFL models on banking data while leaving out a varying number of features with respect to their counterfactual importance rate and comparing against random selection of variables.
@@
[54] tackles an image classification task by calculating Shapley values by SHAP and using them to find the most important pixels for every local FL model and mask the pixels with highest SHAP score. The resulting dataset is used for training the FL model. This method is proposed to protect the FL setup from adversarial attacks, specifically poisoning GAN attacks.
[53] uses random forest (RF) algorithms to detect features causing wrong predictions, with emphasis on detecting malicious attacks on the FL operation. Each participant trains a DL and RF model on its training data. For the samples that are wrongly classified, it calculates the average feature importance of all decision trees with LIME [49] that will result in the same wrong classification. It uses the change in this feature’s importance over time to predict the contribution of each feature in wrongly classifying the data.
@@
[57] implements interpretable adaptive sparse deep networks which exchange NN parameters by means of a multi-level federated network. Whether those weights are shared at the “top sharing level” of the FL architecture depends on the relevance values of the network calculated through layerwise relevance propagation (LRP). The approach provides good diagnostic results even when the FL dataset is under a non-independent identical distribution (NOIID).
@@ A-B Combining FL and explanation methods: applied contributions
[64] trains FL models to segment lung X-ray images and detect signs of pneumonia. Grad-CAM is used to highlight parts of the images that contribute to a detection. It is concluded that a model trained on segmented images has less accuracy, but the pixels highlighted by Grad-CAM focus more on the lung area. It is also reported that training the model in the FL manner helps maintain generalizability and avoid overfitting. A fixed number of FL rounds and greater number of local iterations result into more accuracy.
@@
[41] develops an FL procedure for taxi travel-time prediction based on time-series and geographical data. Authors develop a federated feature attribution aggregation method and test how similar the FL calculated explanations are compared to central calculation. Many XAI techniques are tested, all result in similarly low differences.
[56] compares the use of Shapley values and Lipschitz constant for generating both local and global explanations, and uses this information to update the model. It allows for the personalization of the FL model for each user, so that only the necessary characteristics of the model are retrained, based on the respective needs and the events that it is called to respond.
[89] uses Shapley values computed via SHAP to explain outputs of an FL model trained on edge devices to its operators. It takes the model and the test data as inputs to construct a local linear regression explanation model. Subsequently, the explanatory model computes the Shapley values of classified anomalies and displays them visually. As feature values are measured by sensors, the explanations help operators determine the sensors likely causing an abnormality and make a faster detection response.
[42] proposes to train an FL model to predict the latency in the creation of a network slice. Each slice manager provides data regarding CPU/RAM capacity and usage and serve as FL nodes. The models are evaluated on a local and global level using SHAP, LIME, partial dependent plot (PDP) [50, Ch. 8.1] and RuleFit [51]. Overall, they show that PDP explanations raise privacy concerns since they are run on the client side.
@@
[92] introduces AnoFed, a framework integrating transformer-based Autoencoders (AEs) and Variational Autoencoders (VAEs) with Support Vector Data Description (SVDD) in a federated environment, specifically for ECG anomaly detection, optimizing computational efficiency. A combined design of the VAE/AE and SVDD incorporates kernel density estimation for adaptive anomaly detection. Moreover, it includes an module which explains the anomaly detection output by identifying key segments of the ECG signal that show the maximum reconstruction loss.
[74] presents an ECG-based arrhythmia classification framework which trains convolutional DNNs via FL, and includes an XAI module computing activation mappings in the ECG signal by means of GradCAM. The framework addresses data availability, privacy, and interpretability challenges.
[66] proposes a framework for FL in connected medical devices with blockchain integration for safe model parameter exchange. That framework is showcased by a hybrid implementation of actual hardware nodes and simulated distribution of publicly available datasets in many use cases. Explanations are shown in two cases, however the results and impact of FL are not discussed.
@@ A-C Papers combining FL and interpretable models: methodological contributions
[67] proposes an FL algorithm that builds federated classification models without relying on gradient descent-based methods. Therefore, the class of algorithms that can be learned via FL is not restricted to models whose output is differentiable with respect to the model parameters. Based on AdaBoost [93], it effectively combines gradient-free classifiers which may be learned independently by the FL clients.
[60] trains a fuzzy rule-based system (FRBS) [61] in federation via a one-shot communication scheme where each data silo computes their own FRBS and the individual models are merged by the central server. The proposed FRBS uses a maximum-matching inference rule so the inferred regression function is piecewise linear, which is inherently explainable.
[71] proposes an FL framework for time-series classification , using interpretable, human-understandable time-series features, namely shapelet features, interval features, and dictionary features. The paper claims to guarantee interpretability for the learning-initiating party by ensuring that it can access the aforementioned features without data leakage. To ensure security, the solution incorporates secure feature extraction protocols, secure model training protocols, additive secret sharing schemes, and secure computation protocols.
@@
[76] presents an interpretable data interoperability method for FL called iFedAvg to address the low interoperability due to client data inconsistencies, among other challenges. The iFedAvg method uses personalized layers to adjust for local data shifts, like age differences, directly within input features, which maintains privacy while allowing for direct interpretability. The difference in values of private weight and bias of input layer of each participant captures the inherent shift in data. It was tested on public benchmarks and on a large, real-world Ebola dataset.
[94] introduces a group personalization strategy in FL to address client drift in settings with distinct client partitions. Authors fine-tuned a global FL model with another FL process for each homogeneous client group and then adapted it per client. The method is tested on real-world language modeling datasets and aligns with Bayesian hierarchical modeling principles.
@@
[78] proposes a framework for solving sparse support vector machine (SVM) classification in a distributed fashion. Its performance demonstrated on an electronic health record (EHR) dataset. The proposed algorithm has improved convergence rate compared to several alternatives. Interpretability is assessed by achieving a classifier with fewer features considered as highly important for predictions.
@@ A-D Papers combining FL and interpretable models: applied contributions
[75] implements a FL network based on Gradient Boosting Decision Trees (GBDT). These GBDT are considered transparent models and therefore, their FL network is considered transparent as well. They apply their algorithm on network intrusion data sets and claim that model is interpretable for human experts.
[68] introduces a method in financial risk management for credit scoring and rating with big-data capabilities. Using a VFL framework, it allows multiple agencies to collaboratively train an optimized scorecard model. Performance is showcased on two finance datasets.
@@
[59] investigates client-based attention weight aggregation in a threat-detection task within a cloud scenario, using system logs as input data. Each client predicts local attentions, which are claimed to enhance interpretability, and the central server subsequently aggregates the attention weights to build a saliency map that provides insights on the impact of the different log keys on the threat prediction.
[79] proposes an adaptation for FL of the Cox Proportional Hazards regression model with LASSO regularization. Such an estimator is used as a feature selector in the context of survival analysis and personalized medicine. The inclusion of LASSO regularization enacts a feature selection, contributing to the model’s interpretability.
@@
[95] explores FL of interpretable models in 5G and 6G systems, focusing on automated vehicle networking. The approach addresses gaps in existing AI-based solutions for wireless networks, particularly in vehicle-to-everything (V2X) environments. The methodology offers decentralized, efficient intelligence, enhancing operational trustworthiness and data management.

### v2

@@ V-B FL and Interpretation Methods
A limitation of this study and suggestion for future work is that the FL algorithm (e.g. FedAvg) used in each study was not extracted, which could be a relevant factor in analyzing the impact on explanations. As stated above, adherence to information reporting is important, and future reviews on the field should look closer at this aspect. Further, we conducted our search in 2023. In the ever-changing domain of ML model development, further developments could have been achieved in the meantime.
VI Conclusion
@@
This review studies the interplay between FL and XAI by mapping methodological and experimental contributions. We identified a research gap in which few studies quantify the impact of FL on explanations. Moreover, the impact of FL on model interpretability and explanations remains unclear, revealing the need for studies that quantify such impact. There is a need for rigorous experimental and analytical research to assess how FL training influences the structure of an ML model and its implications for explainability and interpretability. Additional research should also provide guidance for practitioners on the responsible implementation of FL with XAI, particularly in critical application fields such as healthcare, finance, and engineering. This guidance will help ensure the responsible deployment of AI systems that balance model performance with transparency.
@@
Many articles do not use consistent terminology for explainability, which complicates the analysis. It is important for future research to clearly define terms and use standardized nomenclature. Another important finding is that a minority of the studies specify which FL libraries were used. Future publications should include details on the employed libraries or provide source code for their custom implementations. Furthermore, the lack of adequate reporting of data characteristics calls for strict adherence to reporting guidelines to ensure reproducibility, transparency, and auditability. We underscore the need for more structured and transparent research practices at the intersection of FL and XAI. Establishing clear definitions and consistent methodologies will be key in advancing the field and addressing the identified gaps. As demand for FL and XAI continues to grow, particularly in high-stakes application fields such as medicine, the importance of rigorous, transparent, and reproducible research cannot be overstated.
Moreover, while local explanations are frequently applied in FL contexts, no dedicated federated implementations of local explanation methods exist. Aggregating local explanations to obtain global insights often dilutes client-specific patterns, and sharing local explanations may raise privacy concerns. Developing federated-specific local explanation techniques remains an open research challenge.
@@ A-A Combining FL and Explanation Methods: Methodological Contributions
[91] proposes the first counterfactual explanation method for VFL. They show the validity by retraining VFL models on banking data while leaving out a varying number of features concerning their counterfactual importance rate and comparing them against a random selection of variables.
@@
[73] tackles an image classification task by calculating Shapley values by SHAP and using them to find the most important pixels for every local FL model and mask the pixels with the highest SHAP score. The resulting dataset is used for training the FL model. This method is proposed to protect the FL setup from adversarial attacks, specifically poisoning GAN attacks.
[60] uses random forest (RF) algorithms to detect features causing wrong predictions, with an emphasis on detecting malicious attacks on the FL operation. Each participant trains a DL and RF model on its training data. For the samples that are wrongly classified, it calculates the average feature importance of all decision trees with LIME [70] that will result in the same wrong classification. It uses the change in this feature’s importance over time to predict the contribution of each feature in wrongly classifying the data.
@@
[74] implements interpretable adaptive sparse deep networks that exchange NN parameters employing a multi-level federated network. Whether those weights are shared at the “top sharing level” of the FL architecture depends on the relevance values of the network calculated through layerwise relevance propagation (LRP). The approach provides good diagnostic results even when the FL dataset is under a non-independent identical distribution (NOIID).
@@ A-B Combining FL and Explanation Methods: Applied Contributions
[47] trains FL models to segment lung X-ray images and detect signs of pneumonia. Grad-CAM is used to highlight parts of the images that contribute to a detection. It is concluded that a model trained on segmented images has less accuracy, but the pixels highlighted by Grad-CAM focus more on the lung area. It is also reported that training the model in the FL manner helps maintain generalizability and avoid overfitting. A fixed number of FL rounds and a greater number of local iterations result in more accuracy.
@@
[43] develops an FL procedure for taxi travel-time prediction based on time-series and geographical data. Authors develop a federated feature attribution aggregation method and test how similar the FL-calculated explanations are compared to central calculation. Many XAI techniques are tested, and all result in similarly low differences.
[44] compares the use of Shapley values and Lipschitz constant for generating both local and global explanations and uses this information to update the model. It allows for the personalization of the FL model for each user, so that only the necessary characteristics of the model are retrained based on the respective needs and the events it is called to respond to.
[93] uses Shapley values computed via SHAP to explain the outputs of an FL model trained on edge devices to its operators. It takes the model and the test data as inputs to construct a local linear regression explanation model. Subsequently, the explanatory model computes the Shapley values of classified anomalies and displays them visually. As feature values are measured by sensors, the explanations help operators determine the sensors likely causing an abnormality and make a faster detection response.
[63] proposes to train an FL model to predict the latency in the creation of a network slice. Each slice manager provides data regarding CPU/RAM capacity and usage and serves as an FL node. The models are evaluated on a local and global level using SHAP, LIME, partial dependent plot (PDP) [71, Ch. 8.1] and RuleFit [72]. Overall, they show that PDP explanations raise privacy concerns since they are run on the client side.
@@
[96] introduces AnoFed, a framework integrating transformer-based Autoencoders (AEs) and Variational Autoencoders (VAEs) with Support Vector Data Description (SVDD) in a federated environment, specifically for ECG anomaly detection, optimizing computational efficiency. A combined design of the VAE/AE and SVDD incorporates kernel density estimation for adaptive anomaly detection. Moreover, it includes a module that explains the anomaly detection output by identifying key segments of the ECG signal that show the maximum reconstruction loss.
[59] presents an ECG-based arrhythmia classification framework that trains convolutional DNNs via FL and includes an XAI module computing activation mappings in the ECG signal utilizing GradCAM. The framework addresses data availability, privacy, and interpretability challenges.
[50] proposes a framework for FL in connected medical devices with blockchain integration for safe model parameter exchange. That framework is showcased by a hybrid implementation of actual hardware nodes and simulated distribution of publicly available datasets in many use cases. Explanations are shown in two cases, however, the results and impact of FL are not discussed.
@@ A-C Papers combining FL and Interpretable Models: Methodological Contributions
[51] proposes an FL algorithm that builds federated classification models without relying on gradient descent-based methods. Therefore, the class of algorithms that can be learned via FL is not restricted to models whose output is differentiable concerning the model parameters. Based on AdaBoost [97], it effectively combines gradient-free classifiers, which may be learned independently by the FL clients.
[52] trains a fuzzy rule-based system (FRBS) [78] in federation via a one-shot communication scheme where each data silo computes their own FRBS and the individual models are merged by the central server. The proposed FRBS uses a maximum-matching inference rule, so the inferred regression function is piecewise linear, which is inherently explainable.
[56] proposes an FL framework for time-series classification using interpretable, human-understandable time-series features, namely shapelet features, interval features, and dictionary features. The paper claims to guarantee interpretability for the learning-initiating party by ensuring that it can access the aforementioned features without data leakage. To ensure security, the solution incorporates secure feature extraction protocols, secure model training protocols, additive secret sharing schemes, and secure computation protocols.
@@
[62] presents an interpretable data interoperability method for FL called iFedAvg to address the low interoperability due to client data inconsistencies, among other challenges. The iFedAvg method uses personalized layers to adjust for local data shifts, like age differences, directly within input features, which maintains privacy while allowing for direct interpretability. The difference in values of private weight and bias of the input layer of each participant captures the inherent shift in data. It was tested on public benchmarks and a large, real-world Ebola dataset.
[98] introduces a group personalization strategy in FL to address client drift in settings with distinct client partitions. The authors fine-tuned a global FL model with another FL process for each homogeneous client group and then adapted it per client. The method is tested on real-world language modeling datasets and aligns with Bayesian hierarchical modeling principles.
@@
[85] proposes a framework for solving sparse support vector machine (SVM) classification in a distributed fashion. Its performance was demonstrated on an electronic health record (EHR) dataset. The proposed algorithm has an improved convergence rate compared to several alternatives. Interpretability is assessed by achieving a classifier with fewer features considered as highly important for predictions.
@@ A-D Papers Combining FL and Interpretable Models: Applied Contributions
[61] implements an FL network based on Gradient Boosting Decision Trees (GBDT). These GBDT are considered transparent models, and therefore, their FL network is considered transparent as well. They apply their algorithm on network intrusion data sets and claim that the model is interpretable for human experts.
[53] introduces a method in financial risk management for credit scoring and rating with big-data capabilities. Using a VFL framework allows multiple agencies to collaboratively train an optimized scorecard model. Performance is showcased on two finance datasets.
@@
[77] investigates client-based attention weight aggregation in a threat-detection task within a cloud scenario using system logs as input data. Each client predicts local attentions, which are claimed to enhance interpretability, and the central server subsequently aggregates the attention weights to build a saliency map that provides insights on the impact of the different log keys on the threat prediction.
[86] proposes an adaptation for FL of the Cox Proportional Hazards regression model with LASSO regularization. Such an estimator is used as a feature selector in the context of survival analysis and personalized medicine. Including LASSO regularization enacts a feature selection, contributing to the model’s interpretability.
@@
[99] explores the FL of interpretable models in 5G and 6G systems, focusing on automated vehicle networking. The approach addresses gaps in existing AI-based solutions for wireless networks, particularly in vehicle-to-everything (V2X) environments. The methodology offers decentralized, efficient intelligence, enhancing operational trustworthiness and data management.